Ferris Company began January with 4,000 units of its principal product. The cost of each unit is $6. Merchandise transactions for the month of January are as follows: Purchases Date of Purchase Units Unit Cost* Total Cost Jan. 10 3,000 $ 7 $ 21,000 Jan. 18 4,000 8 32,000 Totals 7,000 53,000 * Includes purchase price and cost of freight. Sales Date of Sale Units Jan. 5 2,000 Jan. 12 1,000 Jan. 20 3,000 Total 6,000 5,000 units were on hand at the end of the month. 2. Calculate January's ending inventory and cost of goods sold for the month using LIFO, periodic system.
Answer #1

(1) LIFO (Last-in-First Out) Method:

Under the LIFO method, it is assumed that the last unit of inventory received will be sold first i.e. the last unit received (latest). Hence when we calculate the cost of goods sold & ending inventory value, we will consider the cost of the latest unit first.

Calculation

Date Units Jan. 01 Jan. 05 Opening balance - Sold Balance Purchase - Sold Rate $ per unit 6 6 6 Jan. 10 Jan. 12 Balance 4.000

Cost of Goods Sold (COGS) = A+B+C = 12,000+7,000+24,000 = $ 43,000.

Ending Inventory Value = $ 34,000.

(2) Periodic System:

Under the Periodic Method, we will compute, cost of goods sold on a single period basis. We will use period based on event happen. When we purchase inventory than we will take total inventory remained with us as closing inventory of such period and next cogs of unit sold will be calculated by taking whole closing inventory as a single lot and calculate average cost.

Date Units Rate $ per unit Jan. 01 Jan. 05 Jan. 10 Jan. 12 Opening balance - Sold Balance Purchase Balance -) Sold Balance Pu

Cost of Goods Sold (COGS) = A+B+C = 12,000+6,600+21,900 = $ 40,500.

Ending Inventory Value = $ 36,500.